                                       Source file: src/introduction.tex

INTRODUCTION TO THE ALGEBRAIC STACKS DOCUMENTS

Copyright (c) 2005 Johan de Jong and the algebraic_geometry mailing list.
Permission is granted to copy, distribute and/or modify this document
under the terms of the GNU Free Documentation License, Version 1.2
or any later version published by the Free Software Foundation;
with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts.
A copy of the license is included in the section entitled "GNU
Free Documentation License".

Contents

Section 1. Overview 1 References 1

Section 1. Overview

Besides the book by Laumon and Moret-Bailley, see [LMB00], and the work (in progress) by Fulton et al, we think there is a place for an open source textbook on algebraic stacks. To read about this project and its goals please visit the mailing list. This is also the place to submit patches, etc. To continue reading,

(1) visit the next section: Conventions, Section 1, or

(2) go back to the table of contents: index.html#contents.

References

[LMB00] Gérard Laumon and Laurent Moret-Bailly. Champs algébriques, volume 39 of Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge. Springer-Verlag, 2000.